# Kuro Siwo: 33 billion under the water A global multi-temporal satellite dataset for rapid flood mapping

### v2

@@ Abstract
Global floods, exacerbated by climate change, pose severe threats to human life, infrastructure, and the environment. Recent catastrophic events in Pakistan and New Zealand underscore the urgent need for precise flood mapping to guide restoration efforts, understand vulnerabilities, and prepare for future occurrences. While Synthetic Aperture Radar (SAR) remote sensing offers day-and-night, all-weather imaging capabilities, its application in deep learning for flood segmentation is limited by the lack of large annotated datasets. To address this, we introduce Kuro Siwo, a manually annotated multi-temporal dataset, spanning 43 flood events globally. Our dataset maps more than 338 billion of land, with 33 billion designated as either flooded areas or permanent water bodies. Kuro Siwo includes a highly processed product optimized for flood mapping based on SAR Ground Range Detected, and a primal SAR Single Look Complex product with minimal preprocessing, designed to promote research on the exploitation of both the phase and amplitude information and to offer maximum flexibility for downstream task preprocessing. To leverage advances in large scale self-supervised pretraining methods for remote sensing data, we augment Kuro Siwo with a large unlabeled set of SAR samples. Finally, we provide an extensive benchmark, namely BlackBench, offering strong baselines for a diverse set of flood events from Europe, America, Africa, Asia and Australia. Our code and data will be made available at https://github.com/Orion-AI-Lab/KuroSiwo.
@@ 1 Introduction
The exploitation of satellite data for flood mapping has seen extensive use [10], particularly with the emergence of Sentinel satellite platforms [78]. Specifically, the Sentinel-2 multispectral and Sentinel-1 Synthetic Aperture Radar (SAR) missions offer global, detailed and frequent imaging of the Earthās surface. However, high precipitation and extensive cloud cover during flood events impact multispectral sensors which effectively become blind and unreliable for operational scenarios. In contrast, SAR operates seamlessly day and night, unaffected by weather conditions, making it better suited for operational flood mapping. Thus, our research focuses on developing methods that can rapidly map flooded areas using SAR data.
@@
To overcome this barrier, we curate time series data from Sentinel-1 SAR imagery linked to flood events worldwide, and manually annotate them with the expertise of SAR specialists. We provide two SAR products paired with the resulting reference annotation maps: Ground Range Detected (GRD) SAR with preprocessing tailored for flood mapping and Single Look Complex (SLC) SAR data with minimal preprocessing, containing both phase and amplitude signal. Additionally, we incorporate unlabeled SAR samples, to facilitate the exploration of semi-supervised and large-scale self-supervised learning (SSL). We name the resulting dataset āKuro Siwoā111Named after a poem by sailor-poet Nikos Kavvadias (1910-1975). and release it publicly with the aim to propel research in this field towards developing operational tools that support civil protection authorities for the greater good.
@@ 2.1 Flood mapping datasets
In this section we summarize previous work on the creation of analysis-ready satellite based datasets for flood mapping. Our investigation is focused on three key aspects. First, the provided satellite sensors, second the ground truth generation process and finally the evaluation scheme. Please refer to Tab.Ā 1 for an overview of our analysis.
@@
| UNOSAT [60] | S1 | ā | ā | 10m | 58,128 | 256x256 | 15 | all + 1 new | Post | 2 | 2 | SA |
| SEN12-FLOOD [65] | S1, S2 | ā | ā | 10m | 336 | 512x512 | - | - | Time series | 3 | 2 | CEMS |
| Sen1Floods11 [8] | S1, S2 | ā | ā | 10m | 4,831 | 512x512 | 11 | 1 | Post | 5 | 3 | M, A |
| Ombria [22] | S1, S2 | ā | ā | 10m | 1,688 | 256x256 | 23 | all | Pre, Post | 5 | 2 | CEMS |
| WorldFloods [51] | S2 | ā | ā | 10m | 185,574 | 256x256 | 119 | 6 | Post | 6 | 3 | SA |
| GF-FloodNet [90] | GF2, GF3 | ā | ā | 1.5 - 5m | 13,388 | 256x256 | 8 | all + 3 new | Post | 4 | 2 | SA |
| Global Flood Database [76] | MODIS | ā | ā | 250m | 12,719 | - | 913 | - | Post | 6 | 2 | A |
| S1GFloods [71] | S1 | ā | ā | - | 5,360 | 256x256 | 46 | all + 2 new | Pre, Post | 6 | 2 | SA |
| MMFlood [56] | S1, DEM, OSMh | ā | - | 20m | 1,748 | 2000x2000 | 95 | 34 | Post | 6 | 2 | CEMS |
| RAPID-NRT [88] | S1 | ā | ā | 10m | 559 | - | 4 | - | - | 1 | 2 | A |
| CAU-Flood [31] | S1, S2 | ā | ā | 10m | 18,302 | 256x256 | 18 | 2 | Pre, Post | 4 | 2 | M |
| FloodNet [64] | UAV | ā | - | 1.5cm | 2,343 | 4000x3000 | 1 | 1 | Post | 1 | 9 | SA |
| ETCI2021 [59] | S1 | ā | ā | 20m | 66,810 | 256x256 | 5 | 2 | Post | 3 | 2 | - |
| Kuro Siwo (ours) | S1, DEM | ā | ā | 10m | 67,490 | 224x224 | 43 | 10 | 2 Pre, Post | 6 | 3 | M |
@@
Given the difficulty and cost of satellite imagery photointerpretation, the creation of ground truth masks for large-scale datasets is not trivial, with many works offering automatic or semi-automatic methods to alleviate the need for a laborious labelling process (Tab.Ā 1). Some datasets, like the Global Flood Database and RAPID-NRT, employ automatic annotation methods, introducing noise and inaccuracies in the labels due to their reliance on input data quality. Others, such as UNOSAT, WorldFloods, GF-FloodNet, S1GFloods, and FloodNet, utilize a semi-automatic pipeline where automatically generated labels undergo refinement by human annotators. Itās worth noting that in Sen1Floods11, only samples from a single event were manually annotated, while the rest used a thresholding method without human intervention. Manual labeling through meticulous photointerpretation produces higher-quality ground truth mappings. Nevertheless, this approach is expensive, requiring expert input and potentially introducing annotator bias. Such labels have been produced for the CAU-Flood dataset to generate binary flood/no flood masks.
The proposed evaluation scheme is crucial for assessing the generalization capabilities of any flood mapping method. However, in Sen1Floods11, WorldFloods, CAU-Flood, ETCI2021 and FloodNet, a limited number of events are reserved for testing, potentially providing an insufficient indicator of a modelās performance on unseen conditions. Additionally, in UNOSAT, Ombria, GF-FloodNet, and S1GFloods, samples from all flood events are shuffled and randomly split into training, validation, and test sets, posing a risk of data leakage where samples from the same area and satellite capture may appear in both training and test sets. To address this concern, the authors of UNOSAT, GF-FloodNet, and S1GFloods have also performed evaluation on imagery from a small number of new, unseen flood events, though these events are not included in the published datasets.
@@ 3 Kuro Siwo dataset
Requirements: While substantial efforts have been invested in automated flood mapping, the absence of a diverse and well-curated remote sensing dataset hampers the potential of deep learning methods for this critical task. To address this gap, we present Kuro Siwo, a distinctive flood inundation mapping dataset constructed with specific constraints. Firstly, the dataset is SAR-based, ensuring availability of usable imagery in all weather conditions, day and night, even during production in an operational context. In addition, we stick to SAR imagery from the Sentinel-1A&B satellites, since Copernicus data are available on a free and open basis, and there is strong commitment for the continuity of the mission. Secondly, Kuro Siwo is designed as a multi-temporal dataset to enable distinguishing between permanent water bodies, e.g rivers and lakes, and flooded areas, and also to facilitate testing change detection computer vision architectures. Thirdly, Kuro Siwo aims for diversity, featuring extensive spatio-temporal coverage of major flood events spanning from 2015 to 2022, across six continents and three climate zones. The omission of flood events in Antarctica and polar/cold climate zones reflects their real-life distribution, as these areas typically do not experience floods. Fig.Ā 1 illustrates the spatial coverage of Kuro Siwo, including the distribution of climate zones. Lastly, we prioritize the generation of high-quality flood annotations at a global scale, achieved through manual photointerpretation by a group of SAR experts, surpassing those from existing sources like the Copernicus Emergency Management Service (CEMS).
Input data: For each event, we assemble a triplet of Sentinel-1 data at two forms: a) Level-1 GRD SAR data and b) Level-1 SLC SAR data, both at 10 m spatial resolution. This triplet comprises two pre-event images with varying temporal distances ā eliminating rigid constraints for real-world applications ā and one post-event image acquired as close as possible to the actual event date. The temporal gap depends on the sensor revisit time and the location on earth. We impose that for each flood event, all three SAR images belong to either the descending or ascending imaging geometry to prevent variations in layover, foreshortening and shadow effects [16] within the same sample. Additionally, for each SAR image in Kuro Siwo, both VV (Vertical transmit and Vertical receive) and VH (Vertical transmit and Horizontal receive) polarizations are registered, as studies have demonstrated their complementary value for flood mapping [45, 32].
@@
The above result in problematic annotations that hamper the potential to train robust and accurate deep learning models. For example, the creators of the Ombria [22] and the MMFlood [56] datasets (Tab.Ā 1) that both use CEMS annotations, report low classification accuracies. This was our experience also when we used the original CEMS annotations; our deep learning models achieved less than % for all evaluation metrics. Therefore we decided to invest in updating CEMS annotations through photointerpretation. Fig.Ā 2 exemplifies the annotation improvement attained in Kuro Siwo for a flood event in an agricultural area in France. The errors in CEMS annotations, especially for the permanent waters class, are obvious when carefully examining both polarizations. To discern the disparity between the CEMS and Kuro Siwo annotation for the specific flood event, the two products exhibit Intersection over Union (IoU) of % and % for the permanent water and flood categories, respectively.
@@
Going beyond CEMS: Building on CEMS annotations, which focuses primarily on Europe, results in severe underrepresentation of other continents. Recognizing this, we expand our study on flood events from Asia, Australia, Africa as well as South and North America, aiming for a more balanced spatial distribution.
@@
Dataset split: Scholastically assessing and comparing the capacity of flood mapping models to operate in novel environments is of great importance to reliably evaluate future methods and eventually deploy them in the real world. With that in mind, we construct a challenging evaluation framework, selecting 10 flood events across the globe as testing sites, covering a wide range of environmental conditions representing all six continents and three major climate zones featured in Kuro Siwo. The spatial distribution of training, validation and test flood events in Kuro Siwo is illustrated in Fig.Ā 1 along with the continents representation.
@@
Unlabeled component: In the context of natural hazards and extreme events, instances of positive occurrences, such as floods, volcanic unrest [9], landslides [7] and wildfires [41], are notably rare. This scarcity results in a limited dataset, insufficient for mapping the diverse, intricate and dynamic environmental variables, including water. Furthermore, the number of floods monitored with publicly available SAR imagery is finite, given that Sentinel-1 data are available after 2014 only, while at the same time, acquiring and photointerpreting them all implies a substantial cost. Recognizing these constraints and the importance of generalizing to unseen events, we offer an extensive, unlabeled collection of satellite frame triplets, adhering to the same principles and preprocessing pipeline as the annotated Kuro Siwo set. This resultant dataset encompasses time series with unique SAR samples. Motivated by recent advancements in foundational models for computer vision [40], we release this dataset to encourage the exploration of SSL methods specifically designed for the domain of rapid flood mapping.
@@ 4 BlackBench: An extensive benchmark on rapid flood mapping
Building on Kuro Siwo, we introduce BlackBench, an extensive benchmark employing a large set of powerful models under various configurations to act as strong baselines for future methods. BlackBench includes common semantic segmentation architectures like U-Net [69], DeepLabv3 [12] and UPerNet [86] using both convolutional and transformer backbones like the ResNet [30] family, Swin Transformer [47] and ConvNext [48]. We use two variants of each backbone for UPerNet indicated as Small (S) and Base (B) as defined in [15]. Furthermore, we include a set of models inspired by change detection problems like FC-EF-Diff and FC-EF-Conc [17], which were among the first deep learning models proposed for this task. Additionally, we include SNUNet-CD [24], a densely connected U-Net++ [93] model with a dual-branch encoder, and Changeformer [4], a model consisting of two siamese branches with transformer blocks and a lightweight MLP decoder.
@@
| UNet-ResNet101 | 3 | - | ā | 98.69 | 79.33 | 78.92 | 76.12 | 82.88 |
| DeepLab-ResNet18 | 3 | - | ā | 98.74 | 77.35 | 78.51 | 75.07 | 83.38 |
| DeepLab-ResNet50 | 3 | - | ā | 98.71 | 77.35 | 78.70 | 75.14 | 83.20 |
@@
For each model included in BlackBench, we assess the performance across diverse input scenarios, using GRD data. The input setting varies on two factors. First, in regard to the time series length and second, in regard to the inclusion of elevation information. For all semantic segmentation models we consider time series of either 2 or 3 images varying the number of pre-event captions. For the change detection models we strictly use two captions by selecting only the most recent pre-flood image since this family of architectures typically employs a two-stream input encoder. For each time series length, we examine the performance of the models when we a) include a DEM, b) include slope information (derived from the DEM) and c) provide no elevation information at all. In Tab.Ā 2 we report the performance of the best setting for each architecture. In particular, we report the F1-Score (F1) for each class as well as the overall mean IoU (mIoU). Additionally, we evaluate the F1-Score for the binary task of water detection by combining the predictions for permanent water and flood into one āwaterā class. All semantic segmentation models use pretrained backbones and were trained for 20 epochs. U-Net and DeepLabv3 models use backbones pretrained on ImageNet [18], while UPerNet uses backbones from [15]. Change detection models were trained from scratch for epochs. Finally, we create a baseline model trained via SSL using the unlabeled component of Kuro Siwo, excluding the events from the test set. To model interactions in the temporal dimension as well as make the finetuning process more direct, we treat each Kuro Siwo triplet as one sample. We utilize the Masked Autoencoder [29] (MAE) method employing a vision transformer [21] (ViT) with 24 layers and 16 attention heads as our encoder and train for 100 epochs. We stack the triplets on the channel dimension as done with all segmentation methods in BlackBench. To model the varying temporal distances between the pre-event and the post-event captions, we include three temporal embeddings, i.e. one for each timestep. We encode the year, month and day independently using a sinusoidal encoding and concatenate them. The resulting temporal embeddings are added to the tokens as done with the standard, learnable positional embeddings. For the segmentation downstream task, we add a simple, trainable, convolutional decoder on top of the learnt representations. We evaluate the capacity of our model by a) keeping the encoder frozen and training only the decoder module, and b) finetune the last 12 transformer layers along with the decoder. We will refer to these models as FloodViT and FloodViT-FT respectively. We discuss the insights gained from BlackBench in detail in Sec.Ā 5.

### v3

@@ Abstract
Global flash floods, exacerbated by climate change, pose severe threats to human life, infrastructure, and the environment. Recent catastrophic events in Pakistan and New Zealand underscore the urgent need for precise flood mapping to guide restoration efforts, understand vulnerabilities, and prepare for future occurrences. While Synthetic Aperture Radar (SAR) remote sensing offers day-and-night, all-weather imaging capabilities, its application in deep learning for flood segmentation is limited by the lack of large annotated datasets. To address this, we introduce Kuro Siwo, a manually annotated multi-temporal dataset, spanning 43 flood events globally. Our dataset maps more than 338 billion of land, with 33 billion designated as either flooded areas or permanent water bodies. Kuro Siwo includes a highly processed product optimized for flash flood mapping based on SAR Ground Range Detected, and a primal SAR Single Look Complex product with minimal preprocessing, designed to promote research on the exploitation of both the phase and amplitude information and to offer maximum flexibility for downstream task preprocessing. To leverage advances in large scale self-supervised pretraining methods for remote sensing data, we augment Kuro Siwo with a large unlabeled set of SAR samples. Finally, we provide an extensive benchmark, namely BlackBench, offering strong baselines for a diverse set of flood events globally. All data and code are published in our Github repository: https://github.com/Orion-AI-Lab/KuroSiwo.
@@ 1 Introduction
The exploitation of satellite data for flood mapping has seen extensive use [10], particularly with the emergence of Sentinel satellite platforms [79]. Specifically, the Sentinel-2 multispectral and Sentinel-1 Synthetic Aperture Radar (SAR) missions offer global, detailed and frequent imaging of the Earthβs surface. However, high precipitation and extensive cloud cover during flood events impact multispectral sensors which effectively become blind and unreliable for operational scenarios. In contrast, SAR operates seamlessly day and night, unaffected by weather conditions, making it better suited for operational flood mapping. Thus, our research focuses on developing methods that can rapidly map flooded areas using SAR data.
@@
To overcome this barrier, we curate time series data from Sentinel-1 SAR imagery linked to flood events worldwide, and manually annotate them with the expertise of SAR specialists. We provide two SAR products paired with the resulting reference annotation maps: Ground Range Detected (GRD) SAR with preprocessing tailored for flood mapping and Single Look Complex (SLC) SAR data with minimal preprocessing, containing both phase and amplitude signal. Additionally, we incorporate unlabeled SAR samples, to facilitate the exploration of semi-supervised and large-scale self-supervised learning (SSL). We name the resulting dataset βKuro Siwoβ111Named after a poem by sailor-poet Nikos Kavvadias (1910-1975). and release it publicly with the aim to propel research in this field towards developing operational tools that support civil protection authorities for the greater good.
@@ 2.1 Flood mapping datasets
In this section we summarize previous work on the creation of analysis-ready satellite based datasets for flood mapping. Our investigation is focused on three key aspects. First, the provided satellite sensors, second the ground truth generation process and finally the evaluation scheme. Please refer to Tab.Β 1 for an overview of our analysis.
@@
UNOSAT [60] | S1 | β | β | 10m | 58,128 | 256x256 | 15 | all + 1 new | Post | 2 | 2 | SA |
SEN12-FLOOD [65] | S1, S2 | β | β | 10m | 336 | 512x512 | - | - | Time series | 3 | 2 | CEMS |
Sen1Floods11 [8] | S1, S2 | β | β | 10m | 4,831 | 512x512 | 11 | 1 | Post | 5 | 3 | M, A |
Ombria [22] | S1, S2 | β | β | 10m | 1,688 | 256x256 | 23 | all | Pre, Post | 5 | 2 | CEMS |
WorldFloods [51] | S2 | β | β | 10m | 185,574 | 256x256 | 119 | 6 | Post | 6 | 3 | SA |
GF-FloodNet [91] | GF2, GF3 | β | β | 1.5 - 5m | 13,388 | 256x256 | 8 | all + 3 new | Post | 4 | 2 | SA |
Global Flood Database [77] | MODIS | β | β | 250m | 12,719 | - | 913 | - | Post | 6 | 2 | A |
S1GFloods [71] | S1 | β | β | - | 5,360 | 256x256 | 46 | all + 2 new | Pre, Post | 6 | 2 | SA |
MMFlood [56] | S1, DEM, OSMh | β | - | 20m | 1,748 | 2000x2000 | 95 | 34 | Post | 6 | 2 | CEMS |
RAPID-NRT [89] | S1 | β | β | 10m | 559 | - | 4 | - | - | 1 | 2 | A |
CAU-Flood [31] | S1, S2 | β | β | 10m | 18,302 | 256x256 | 18 | 2 | Pre, Post | 4 | 2 | M |
FloodNet [64] | UAV | β | - | 1.5cm | 2,343 | 4000x3000 | 1 | 1 | Post | 1 | 9 | SA |
ETCI2021 [59] | S1 | β | β | 20m | 66,810 | 256x256 | 5 | 2 | Post | 3 | 2 | - |
Kuro Siwo (labelled) | S1, DEM | β | β | 10m | 67,490 | 224x224 | 43 | 10 | 2 Pre, Post | 6 | 3 | M |
Kuro Siwo (unlabelled) | S1, DEM | β | β | 10m | 466,357 | 224x224 | 43 | 10 | 2 Pre, Post | 6 | 3 | M |
@@
Given the difficulty and cost of satellite imagery photointerpretation, the creation of ground truth masks for large-scale datasets is not trivial, with many works offering automatic or semi-automatic methods to alleviate the need for a laborious labelling process (Tab.Β 1). Some datasets, like the Global Flood Database and RAPID-NRT, employ automatic annotation methods, introducing noise and inaccuracies in the labels due to their reliance on input data quality. Others, such as UNOSAT, WorldFloods, GF-FloodNet, S1GFloods, and FloodNet, utilize a semi-automatic pipeline where automatically generated labels undergo refinement by human annotators. Itβs worth noting that in Sen1Floods11, only samples from a single event were manually annotated, while the rest used a thresholding method without human intervention. Manual labeling through meticulous photointerpretation produces higher-quality ground truth mappings. Nevertheless, this approach is expensive, requiring expert input and potentially introducing annotator bias. Such labels have been produced for the CAU-Flood dataset to generate binary flood/no flood masks.
The proposed evaluation scheme is crucial for assessing the generalization capabilities of any flood mapping method. However, in Sen1Floods11, WorldFloods, CAU-Flood, ETCI2021 and FloodNet, a limited number of events are reserved for testing, potentially providing an insufficient indicator of a modelβs performance on unseen conditions. Additionally, in UNOSAT, Ombria, GF-FloodNet, and S1GFloods, samples from all flood events are shuffled and randomly split into training, validation, and test sets, posing a risk of data leakage where samples from the same area and satellite capture may appear in both training and test sets. To address this concern, the authors of UNOSAT, GF-FloodNet, and S1GFloods have also performed evaluation on imagery from a small number of new, unseen flood events, though these events are not included in the published datasets.
@@ 3 Kuro Siwo dataset
Requirements: While substantial efforts have been invested in automated flood mapping, the absence of a diverse and well-curated remote sensing dataset hampers the potential of deep learning methods for this critical task. To address this gap, we present Kuro Siwo, a distinctive flood inundation mapping dataset constructed with specific constraints. Firstly, the dataset is SAR-based, ensuring availability of usable imagery in all weather conditions, day and night, even during production in an operational context. In addition, we stick to SAR imagery from the Sentinel-1A&B satellites, since Copernicus data are available on a free and open basis, and there is strong commitment for the continuity of the mission. Secondly, Kuro Siwo is designed as a multi-temporal dataset to enable distinguishing between permanent water bodies, e.g rivers and lakes, and flooded areas, and also to facilitate testing change detection computer vision architectures. Thirdly, Kuro Siwo aims for diversity, featuring extensive spatio-temporal coverage of major flood events spanning from 2015 to 2022, across six continents and three climate zones. The omission of flood events in Antarctica and polar/cold climate zones reflects their real-life distribution, as these areas typically do not experience floods. Fig.Β 1 illustrates the spatial coverage of Kuro Siwo, including the distribution of climate zones. Lastly, we prioritize the generation of high-quality flood annotations at a global scale, achieved through manual photointerpretation by a group of SAR experts, surpassing those from existing sources like the Copernicus Emergency Management Service (CEMS).
Input data: For each event, we assemble a triplet of Sentinel-1 data at two forms: a) Level-1 GRD SAR data and b) Level-1 SLC SAR data, both at 10 m spatial resolution. This triplet comprises two pre-event images with varying temporal distances β eliminating rigid constraints for real-world applications β and one post-event image acquired as close as possible to the actual event date. The temporal gap depends on the sensor revisit time and the location on earth, resulting in a mean of 3.6 days (with std of 6.07 days) and a median of 1 day in post-event captions. We impose that for each flood event, all three SAR images belong to either the descending or ascending imaging geometry to prevent variations in layover, foreshortening and shadow effects [16] within the same sample. Additionally, for each SAR image in Kuro Siwo, both VV (Vertical transmit and Vertical receive) and VH (Vertical transmit and Horizontal receive) polarizations are registered, as studies have demonstrated their complementary value for flood mapping [45, 32].
@@
The above result in problematic annotations that hamper the potential to train robust and accurate deep learning models. For example, the creators of the Ombria [22] and the MMFlood [56] datasets (Tab.Β 1) that both use CEMS annotations, report low classification accuracies. This was our experience also when we used the original CEMS annotations; our deep learning models achieved less than % for all evaluation metrics. Therefore we decided to invest in updating CEMS annotations through photointerpretation. Fig.Β 2 exemplifies the annotation improvement attained in Kuro Siwo for a flood event in an agricultural area in France. The errors in CEMS annotations, especially for the permanent waters class, are obvious when carefully examining both polarizations. To discern the disparity between the CEMS and Kuro Siwo annotation for the specific flood event, the two products exhibit Intersection over Union (IoU) of % and % for the permanent water and flood categories, respectively.
@@
Going beyond CEMS: Building on CEMS annotations, which focuses primarily on Europe, results in severe underrepresentation of other continents. Recognizing this, we expand our study on flood events from Asia, Australia, Africa as well as South and North America, aiming for a more balanced spatial distribution. These events cover extensive areas providing a substantial number of additional training patches. Floods in Europe are fragmented small scale events, as opposed to large flood events encountered in other continents, e.g. Asia. Consequently, while most events in Kuro Siwo are from Europe compared to Asia, the latter contributes with more samples.
@@
Dataset split: Scholastically assessing and comparing the capacity of flood mapping models to operate in novel environments is of great importance to reliably evaluate future methods and eventually deploy them in the real world. With that in mind, we construct a challenging evaluation framework, selecting 10 flood events across the globe as testing sites, covering a wide range of environmental conditions representing all six continents and three major climate zones featured in Kuro Siwo. The spatial distribution of training, validation and test flood events in Kuro Siwo is illustrated in Fig.Β 1 along with the continents representation.
@@
Unlabeled component: In the context of natural hazards and extreme events, instances of positive occurrences, such as floods, volcanic unrest [9], landslides [7] and wildfires [41], are notably rare. This scarcity results in a limited dataset, insufficient for mapping the diverse, intricate and dynamic environmental variables, including water. Furthermore, the number of flash floods monitored with publicly available SAR imagery is finite, given that Sentinel-1 data are available after 2014 only, while at the same time, acquiring and photointerpreting them all implies a substantial cost. Recognizing these constraints and the importance of generalizing to unseen events, we offer an extensive, unlabeled collection of satellite frame triplets, adhering to the same principles and preprocessing pipeline as the annotated Kuro Siwo set. This resultant dataset encompasses time series with unique SAR samples. Motivated by recent advancements in foundational models for computer vision [40], we release this dataset to encourage the exploration of SSL methods specifically designed for the domain of rapid flood mapping.
@@ 4 BlackBench: An extensive benchmark on rapid flood mapping
Building on Kuro Siwo, we introduce BlackBench, an extensive benchmark employing a large set of powerful models under various configurations to act as strong baselines for future methods. BlackBench includes common semantic segmentation architectures like U-Net [69], DeepLabv3 [12] and UPerNet [87] using both convolutional and transformer backbones like the ResNet [30] family, Swin Transformer [47] and ConvNext [48]. We use two variants of each backbone for UPerNet indicated as Small (S) and Base (B) as defined in [15]. Furthermore, we include a set of models inspired by change detection problems, like FC-EF-Diff and FC-EF-Conc [17], which were among the first deep learning models proposed for this task. Additionally, we include SNUNet-CD [24], a densely connected U-Net++ [94] model with a dual-branch encoder, and Changeformer [4], a model consisting of two siamese branches with transformer blocks and a lightweight MLP decoder. Finally, we explore ConvLSTM [73], a recurrent convolutional architecture suitable for time-series data. We employ an encoder-decoder architecture, and we select only the last output map for an N-to-1 scheme.
@@
UNet-ResNet101 | 3 | - | β | 98.69 | 79.33 | 78.92 | 76.12 | 82.88 |
DeepLab-ResNet18 | 3 | - | β | 98.74 | 77.35 | 78.51 | 75.07 | 83.38 |
DeepLab-ResNet50 | 3 | - | β | 98.71 | 77.35 | 78.70 | 75.14 | 83.20 |
@@
ConvLSTM | 3 | - | - | 98.66 | 76.57 | 77.53 | 74.23 | 81.96 |
For each model included in BlackBench, we assess the performance across diverse input scenarios, using GRD data. The input setting varies on two factors. First, in regard to the time series length and second, in regard to the inclusion of elevation information. For all semantic segmentation models we consider time series of either 2 or 3 images varying the number of pre-event captions. For the change detection models we strictly use two captions by selecting only the most recent pre-flood image since this family of architectures typically employs a two-stream input encoder. For each time series length, we examine the performance of the models when we a) include a DEM, b) include slope information (derived from the DEM) and c) provide no elevation information at all. In Tab.Β 2 we report the performance of the best setting for each architecture. In particular, we report the F1-Score (F1) for each class as well as the overall mean IoU (mIoU). Additionally, we evaluate the F1-Score for the binary task of water detection by combining the predictions for permanent water and flood into one βwaterβ class. All semantic segmentation models use pretrained backbones and were trained for 20 epochs. U-Net and DeepLabv3 models use backbones pretrained on ImageNet [18], while UPerNet uses backbones from [15]. Change detection and recurrent models were trained from scratch for epochs. Finally, we create a baseline model trained via SSL using the unlabeled component of Kuro Siwo, excluding the events from the test set. To model interactions in the temporal dimension as well as make the finetuning process more direct, we treat each Kuro Siwo triplet as one sample. We utilize the Masked Autoencoder [29] (MAE) method employing a vision transformer [21] (ViT) with 24 layers and 16 attention heads as our encoder and train for 100 epochs. We stack the triplets on the channel dimension as done with all segmentation methods in BlackBench. To model the varying temporal distances between the pre-event and the post-event captions, we include three temporal embeddings, i.e. one for each timestep. We encode the year, month and day independently using a sinusoidal encoding and concatenate them. The resulting temporal embeddings are added to the tokens as done with the standard, learnable positional embeddings. For the segmentation downstream task, we add a simple, trainable, convolutional decoder on top of the learnt representations. We evaluate the capacity of our model by a) keeping the encoder frozen and training only the decoder module, and b) finetune the last 12 transformer layers along with the decoder. We will refer to these models as FloodViT and FloodViT-FT respectively. We discuss the insights gained from BlackBench in detail in Sec.Β 5. Furthermore, we provide BlackBench results for the SLC component in the Supplementary material.